# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 2.2.3. Decoding Strategies
2.3. Obtain SLM from LLM
Obtaining a small language model (SLM) from a large language model (LLM) is crucial for deploying in resource-constrained environments. Instead of training from scratch, leveraging an LLM allows for knowledge transfer, enabling SLMs to retain much of the LLM’s linguistic and domain knowledge with reduced training time and data. To obtain SLMs from LLMs, three primary techniques are used: pruning, knowledge distillation, and quantization. Pruning removes less critical parameters, reducing model size while aiming to maintain performance. Knowledge distillation transfers knowledge from a large teacher model to a smaller student model, preserving much of the original model’s understanding. Quantization decreases parameter precision, significantly lowering memory and computation needs with minimal impact on accuracy. These methods balance size reduction, efficiency, and performance retention.
Figure 5. Visualizations of unstructured and structured pruning.
@@ 2.3.1. Pruning
Pruning is a technique used to reduce a model’s size and computational requirements (e.g., LLMs) without significantly sacrificing its performance (Han et al., 2015). This process involves identifying and removing less important or redundant parameters and components from the model. The primary goal of LLM pruning is to make the model more efficient, faster, and suitable for deployment in resource-constrained environments. Typically, pruning can be categorized into two main types: unstructured pruning and structured pruning (Zhu et al., 2023a; Wang et al., 2024a). An illustration of unstructured pruning and structured pruning is shown in Figure 5.
@@
Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023e; Shao et al., 2024; Das et al., 2023) prunes an LLM by removing weights individually without considering its internal structure. The least significant parameters are pruned according to specific criteria (e.g. magnitude or impact on the output). This method can achieve significant compression while maintaining performance. However, it can also lead to irregular memory access patterns and reduced hardware efficiency because the pruned model lacks a regular structure. SparseGPT (Frantar and Alistarh, 2023) is a representative unstructured pruning method that can reduce large-scale GPT models like OPT-175B (Zhang et al., 2022) and BLOOM-176B (Le Scao et al., 2023) to up to 60% sparsity using a novel sparse regression solver. Wanda (Sun et al., 2024c) combines weight magnitudes with input activations to efficiently identify and discard less impactful parameters. It operates in a single forward pass, rapidly achieving high sparsity without retraining. It is also worth noting that recent studies specifically address the compatibility issues between pruning and Low-rank Adaptation (LoRA) (Hu et al., 2021), such as LoRAPrune (Zhang et al., 2023a).
@@
Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024c; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024g; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b), which prunes an LLM by targeting entire structural components—such as neurons, channels, or layers—rather. This approach allows for a direct reduction in dimensionality, thus efficiently reducing model complexity and memory usage. Although structured pruning may lead to higher accuracy degradation than unstructured pruning, it simplifies implementation without requiring specialized hardware. ShortGPT (Men et al., 2024) proposes the Block Influence (BI) metric, which measures the significance of each layer based on its transformation of hidden states. Essentially, a transformer block’s influence is measured by how much it alters the hidden states. By calculating BI scores, ShortGPT determines which layers contribute minimally to the overall performance and removes these low-importance layers. This simple yet effective layer removal strategy significantly reduces the model’s parameters and computational requirements. LLM Pruner (Ma et al., 2023c) offers a method to efficiently prune LLMs without access to the original training dataset. It employs a three-step compression pipeline: Discovery (identifying interdependent structures), Estimation (evaluating the performance impact of each group), and Recovery (post-training to address performance loss). NutePrune (Li et al., 2024c) enhances structured pruning with a Numerous-teacher method, employing variable sparsity masks and LoRA modules to guide the pruning process. This approach effectively reduces model size and complexity.
@@ 2.3.2. Knowledge Distillation
Knowledge distillation (KD) compresses a larger teacher model into a smaller student model by training the student to mimic the teacher’s outputs (Hinton et al., 2015). This enables the student to retain much of the teacher’s capabilities with fewer parameters, making it ideal for scaling down LLMs for resource-limited environments while maintaining performance. KD can be categorized into white-box and black-box approaches (Wang et al., 2024a; Zhu et al., 2023a; Yang et al., 2024c) as shown in Figure 6. In White-Box KD, the student has access to the teacher’s internal states or output distributions (Zhang et al., 2023b; Gu et al., 2024; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024). Generalized Knowledge Distillation (GKD) (Ko et al., 2024) introduces skew KL divergence to stabilize gradients and enhance performance, using an adaptive off-policy approach to minimize noisy feedback and improve efficiency. Black-Box KD relies only on teacher outputs without having access to model internals (Peng et al., 2023a; Chen et al., 2023d; Wang et al., 2023f). Methods like Distilling Step-by-Step (Hsieh et al., 2023) use teacher-generated rationales to train smaller models, improving performance with fewer examples. LaMini-LM (Wu et al., 2024b) creates a diverse instruction dataset with GPT-3.5 Turbo responses, enabling robust performance in smaller models.
@@
Mehtods Bit Type Technical Contribution Problems
@@
OneBit (Xu et al., 2024) 1-bit PTQ Quantization-aware knowledge distillation 1-bit quantization
@@
2.3.3. Quantization
Quantization reduces the storage and computational demands of LLMs by converting floating-point representations into lower-precision formats, significantly cutting both storage requirements and computational complexity. Existing methods fall into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). Figure 7 illustrates the two quantization methods. Post-Training Quantization, applied after training, simplifies model compression without altering the architecture or requiring retraining, though it may result in precision loss. Consider a group or block of weights $\mathbf{w}$; the linear operation can be expressed as $y=\mathbf{w}\mathbf{x}$, while the quantized version is given by $y=Q(\mathbf{w})\mathbf{x}$. Generally, the quantization function $Q$ is defined as (Lin et al., 2024c):
$Q(\mathbf{w})=\Delta\cdot\text{Round}\left(\frac{\mathbf{w}}{\Delta}\right),\quad\Delta=\frac{\text{max}(|\mathbf{w}|)}{2^{N-1}},$
where $N$ is the number of quantization bits, and $\Delta$ is the quantization scale factor determined by the absolute maximum value of $\mathbf{w}$. Quantization-Aware Training (QAT) enhances LLM efficiency by incorporating quantization directly into the training process, often resulting in higher accuracy compared to PTQ. During QAT, the forward pass utilizes quantized weights $Q(\mathbf{W})$ and activations $Q(\mathbf{X})$, while retaining full-precision values during the backward pass and for updating gradients to ensure stable learning dynamics. The comparisons of the post-training quantization methods are summarized in Table 1, detailing precision, addressed problems, and technical contributions of each method.
@@ 2.3.4. Low-Rank Techniques
Low-rank techniques compress LLMs by approximating high-dimensional weight matrices with two lower-dimensional matrices, reducing computational and memory requirements. A matrix $\mathbf{W}$ of size $m\times n$ is approximated as $\mathbf{W}\approx\mathbf{A}\times\mathbf{B}$, where $\mathbf{A}$ is $m\times r$ and $\mathbf{B}$ is $r\times n$, with $r$ much smaller than $m$ or $n$, significantly reducing the number of parameters. Building on this concept, Ji et al. (2024) propose a low-rank compression method tailored for LLMs, leveraging the observation that while LLMs have high-rank weights, their feature interactions tend to exhibit low-rank properties. The method estimates feature distributions using pooled covariance matrices and allocates distinct compression ratios to layers based on their sensitivity to low-rank compression. A Bayesian optimization strategy, using a Gaussian process as the surrogate model, optimizes the allocation of low-rank dimensions, ensuring the model maintains performance while achieving significant compression. Transitioning from model compression to fine-tuning, Cho et al. (2024) tackles system and data heterogeneity with the HETLORA method, which uses heterogeneous low-rank approximations to accommodate the diverse capabilities of clients and data complexities. By combining local rank self-pruning with sparsity-weighted aggregation, it balances high and low-rank LoRA modules, improving convergence speed and performance compared to uniform approaches.
2.4. Comparison between Different Compression Methods
Table 2. Comparison of Model Compression Techniques
Criteria Pruning Knowledge Distillation Quantization Low-Rank Techniques
Definition Removes unneeded parameters Transfers knowledge from a larger to a smaller model Lowers the precision of parameters Uses low-rank decomposition on weights
Goal Reduces size and computation Shrinks model while retaining performance Decreases size and speeds up processing Reduces parameters and computation
Method Cuts weights or layers based on importance Smaller model mimics larger model’s output Converts parameters to lower precision Decomposes matrices into smaller components
Advantages Reduces size and computation significantly Preserves performance in smaller models Speeds up inference, less storage Efficient, mostly preserves performance
Disadvantages May reduce accuracy, irregular memory access Resource-intensive, requires large teacher model Potential accuracy loss, may need specific hardware Effectiveness varies, requires rank selection
Model Size Impact High reduction Significant reduction through knowledge transfer High, tied to precision reduction Moderate, reduces redundant parameters
Performance Impact Possible degradation if over-pruned Maintains if well distilled Minor to moderate loss, depends on method Mostly retains performance, may need tuning
Complexity Moderate, needs parameter evaluation High, involves dual-model training Moderate, varies with precision level Moderate, requires matrix factorization
Use Cases Resource-limited settings Efficient model creation for limited resources Fast inference needs, edge devices When weight matrices are redundant
To summarize the key differences between pruning, knowledge distillation, quantization, and low-rank techniques for model compression, we present a comparative table highlighting their definitions, goals, advantages, disadvantages, and typical use cases, as shown in Table 2.
@@ 3. Advanced Enhancement Strategies for Small Language Models
Training from Scratch MindLLM (Yang et al., 2023a) Bilingual pre-trained lightweight models with advanced architectural features
MobiLlama (Thawakar et al., 2024) An on-device SLM with dual objectives of model capability and efficiency through parameter sharing
MobileLLM (Liu et al., 2024h) A deep and thin architecture with advanced techniques to optimize LLM deployment on mobile devices
Supervised Fine-tuning MobileBERT (Sun et al., 2020) Compact BERT that can fine-tune on downstream tasks.
Alpaca 7B (Taori et al., 2023) 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a).
RLHF (Ouyang et al., 2022) Gather human-preferred data, train a reward model, and fine-tune the LM using reinforcement learning.
DPO (Rafailov et al., 2024) Adjust the log probabilities of preferred versus non-preferred responses using a dynamic weighting mechanism, preventing model degradation issues.
Data Quality in KD TinyStory (Eldan and Li, 2023) Small language models can generate coherent stories using a child-friendly dataset and foundational vocabulary
AS-ES (Xi et al., 2024) Categorize reasoning steps into extractive and abstractive segments for enhanced small model CoT capabilities
Self-Amplify (Bhan et al., 2024) A proxy-free, self-generating rationale method for small language models to automate CoT data annotation
Distillation for SLM GKD (Agarwal et al., 2024) Align training and inference distributions using on-policy sequences and flexible divergence measures
DistiLLM (Ko et al., 2024) A skew KL divergence loss for stability and an adaptive off-policy approach for efficient student-generated output utilization.
Adapt-and-Distill (Yao et al., 2021) Enhances small models by first domain adapting both teacher and student models before distillation
Quantization SmoothQuant (Xiao et al., 2023) 8-bit quantization by balancing quantization difficulty between activations and weights using a per-channel scaling transformation
BiLLM (Huang et al., 2024b) Post-training quantization by using a Hessian-based metric to identify salient weights and applying binary residual approximation and optimal splitting search for precise quantization.
LLM-QAT (Liu et al., 2023a) Quantization by data-free knowledge distillation and fine-tuning with logit distillation from the full-precision model
PB-LLM (Shang et al., 2023) Selectively binarizes non-salient weights while preserving salient ones in higher precision, balancing model compression and accuracy
OneBit (Xu et al., 2024) Sign-Value-Independent Decomposition to achieve near 1-bit quantization, balancing extreme compression with minimal performance loss
BitNet (Wang et al., 2023e) A 1-bit Transformer architecture with BitLinear layers and quantization-aware training, optimizing low-precision representation for better accuracy and training stability.
BitNet b1.58 (Ma et al., 2024) Enhanced BitNet by introducing a ternary weight system, achieving full-precision performance from 3 billion parameters with reduced memory and latency costs
LLMs for
SLM
Ma et al. (Ma et al., 2023a) An adaptive filtering and re-ranking paradigm combining LLMs and SLMs improves Information Extraction tasks
MoQE (Kim et al., 2023a) Apply quantization only to expert weights and achieve better performance than the dense model trained on the same dataset.
SLM-RAG (Liu et al., 2024g) Suggests that SLMs equipped with Retrieval-Augmented Generation (RAG) can perform comparably to Large Language Models (LLMs).
@@ 3.1. Innovative Training Methods for Small Language Models from Scratch
Architecture Design for SLMs When designing SLM architectures, parameter-sharing techniques are employed to minimize space usage and reduce the model’s size. As shown in the first part of Figure 8, parameter sharing is achieved by two approaches: (i) a single Feed-Forward Network (FFN) module is shared by every transformer layer, and (ii) entire transformer blocks are shared. For example, to design language models for scenarios that require on-device processing ability and energy, FFN layer sharing/reusing (see Figure 8 (1)) can be a potential option. By sharing FFN layers, the model can maintain a smaller size while still benefiting from the depth and complexity gained through repeated processing of input data. This technique is firstly applied in MobiLlama (Thawakar et al., 2024) which surpasses the performance of existing SLMs of comparable size. Generally, deeper and thinner models consistently perform better than shallower and wider ones (Liu et al., 2024h). Based on this observation, Transformer Block-wise Sharing is another parameter-sharing approach that maintains depth and complexity. There are different transformer block-wise sharing strategies such as repeating the transformer blocks all over again or repeating the immediate transformer block. Experiments show that simply reusing the same transformer blocks can still improve performance (Liu et al., 2024h). This improvement likely comes from the fact that repeating the blocks effectively maintains the model’s depth and complexity, allowing it to capture more information without adding new parameters. While repeat-all-over sharing generally yields better performance, block-wise immediate sharing optimizes cache utilization. This is because the shared weights remain in the cache and can be processed again immediately. This technique is applied in MobileLLMs (Liu et al., 2024h) which has 125M and 350M parameters. MobileLLMs demonstrate performance improvements of 2.7% and 4.3%, respectively, compared to previous models with equivalent parameters. Moreover, they exhibit accuracy comparable to LLaMa-2-7B on API call tasks, highlighting the capabilities of smaller models in mobile environments.
Data Construction For small language models, the emphasis on data quality surpasses that of quantity and diversity (Yang et al., 2023a). Experiments demonstrate that using a quality filtering approach to remove low-quality data can lead to improved performance in SLMs (Yang et al., 2023a). Unlike large models, which can handle diverse and large datasets, SLMs benefit more from cleaner, high-quality data probably due to their limited capacity against noise. Generally, data processing has several steps: (i) Format cleaning, which involves removing HTML, CSS, and JS identifiers, as well as non-textual elements, to ensure clean text input; (ii) Low-quality content filtering that excludes webpages with low text-to-content ratios; (iii) Deduplication using Locality-Sensitive Hashing (LSH), particularly the SimHash algorithm (Dasgupta et al., 2011; Sadowski and Levin, 2007), to prevent content redundancy; (iv) Sensitive information exclusion through heuristics and a sensitive word vocabulary to filter out offensive and illegal content, with special tokens replacing private information; and (v) A self-repeating content filter that eliminates repetitive phrases typical of advertisements, enhancing the training dataset’s information value (Chen et al., 2024b; Yang et al., 2023a). These steps collectively ensure that training data has high-quality, informative texts. SLMs also significantly benefit from these techniques. For example, MindLLMs (Yang et al., 2023a), which are bilingual lightweight language models (available in 1.3B and 3B versions), adopt these data processing techniques and achieve improved capability acquisition, while mitigating issues like catastrophic forgetting.
@@
Insights: We draw several key insights from the training techniques of SLMs: For parameter sharing techniques, maintaining complexity and depth of model structure is essential for maintaining model performance (e.g., shared Feed-Forward Networks (Thawakar et al., 2024) and transformer blocks) (Liu et al., 2024h). Data quality is more important than data quantity in the effectiveness of SLMs (Yang et al., 2023a). Different from LLMs, we can take advantage of the compact nature of SLMs and employ more flexible training strategies, such as multiple-round training (Tang et al., 2024a).

### v2

@@ 2.2.3. Decoding Strategies
2.3. Obtain SLMs from LLMs
Obtaining an SLM from an LLM is crucial for deploying in resource-constrained environments. Instead of training from scratch, leveraging an LLM allows for knowledge transfer, enabling SLMs to retain much of the LLM’s linguistic and domain knowledge with reduced training time and data. To obtain SLMs from LLMs, three primary techniques are used: pruning, knowledge distillation, and quantization. Pruning removes less critical parameters, reducing model size while aiming to maintain performance. Knowledge distillation transfers knowledge from a large teacher model to a smaller student model, preserving much of the original model’s understanding. Quantization decreases parameter precision, significantly lowering memory and computation needs with minimal impact on accuracy. These methods balance size reduction, efficiency, and performance retention.
Figure 7. Unstructured and structured pruning.
@@ 2.3.1. Pruning
Pruning is a technique used to reduce a model’s size and computational requirements (e.g., LLMs) without significantly sacrificing its performance (Han et al., 2015). This process involves identifying and removing less important or redundant parameters and components from the model. The primary goal of LLM pruning is to make the model more efficient, faster, and suitable for deployment in resource-constrained environments. Typically, pruning can be categorized into two main types: unstructured pruning and structured pruning. An illustration of unstructured pruning and structured pruning is shown in Figure 7.
@@
Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023f; Shao et al., 2024; Das et al., 2023) prunes an LLM by removing weights individually without considering its internal structure. The least significant parameters are pruned according to specific criteria (e.g. magnitude or impact on the output). This method can achieve significant compression while maintaining performance. However, it can also lead to irregular memory access patterns and reduced hardware efficiency because the pruned model lacks a regular structure. SparseGPT (Frantar and Alistarh, 2023) is a representative unstructured pruning method that can reduce large-scale GPT models like OPT-175B (Zhang et al., 2022) and BLOOM-176B (Le Scao et al., 2023) to up to 60% sparsity using a novel sparse regression solver. Wanda (Sun et al., 2024c) combines weight magnitudes with input activations to efficiently identify and discard less impactful parameters. It operates in a single forward pass, rapidly achieving high sparsity without retraining. It is also worth noting that recent studies specifically address the compatibility issues between pruning and Low-rank Adaptation (LoRA) (Hu et al., 2021), such as LoRAPrune (Zhang et al., 2023a).
@@
Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024d; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024h; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b; Shen et al., 2024b, 2022; Gao et al., 2024), which prunes an LLM by targeting entire structural components—such as neurons, channels, or layers—rather. This approach allows for a direct reduction in dimensionality, thus efficiently reducing model complexity and memory usage. Although structured pruning may lead to higher accuracy degradation than unstructured pruning, it simplifies implementation without requiring specialized hardware. ShortGPT (Men et al., 2024) proposes the Block Influence (BI) metric, which measures the significance of each layer based on its transformation of hidden states. Essentially, a transformer block’s influence is measured by how much it alters the hidden states. By calculating BI scores, ShortGPT determines which layers contribute minimally to the overall performance and removes these low-importance layers. This simple yet effective layer removal strategy significantly reduces the model’s parameters and computational requirements. LLM Pruner (Ma et al., 2023c) offers a method to efficiently prune LLMs without access to the original training dataset. It employs a three-step compression pipeline: Discovery (identifying interdependent structures), Estimation (evaluating the performance impact of each group), and Recovery (post-training to address performance loss). NutePrune (Li et al., 2024d) enhances structured pruning with a Numerous-teacher method, employing variable sparsity masks and LoRA modules to guide the pruning process. This approach effectively reduces model size and complexity. COST-EFF (Shen et al., 2022) introduces a slenderized backbone—a form of structured pruning—and a multi-exit model that employs task-specific calibration through knowledge distillation. This slenderization reduces the model’s spatial footprint, while the multi-exit strategy effectively balances utility and runtime costs. To enhance the flexibility of structural pruning, DISP-LLM (Gao et al., 2024) breaks the structural dependencies in regular methods by allowing different layers to have different subsets of features along the embedding dimension.
@@ 2.3.2. Knowledge Distillation
Knowledge distillation (KD) compresses a larger teacher model into a smaller student model by training the student to mimic the teacher’s outputs (Hinton et al., 2015). This enables the student to retain much of the teacher’s capabilities with fewer parameters, making it ideal for scaling down LLMs for resource-limited environments while maintaining performance. KD can be categorized into white-box and black-box approaches (Wang et al., 2024c; Zhu et al., 2023a; Yang et al., 2024c) as shown in Figure 8. In White-Box KD, the student has access to the teacher’s internal states or output distributions (Zhang et al., 2023b; Gu et al., 2024a; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024; Wu et al., 2024b). Generalized Knowledge Distillation (GKD) (Ko et al., 2024) introduces skew KL divergence to stabilize gradients and enhance performance, using an adaptive off-policy approach to minimize noisy feedback and improve efficiency. Black-Box KD relies only on teacher outputs without having access to model internals (Peng et al., 2023b; Chen et al., 2023c; Wang et al., 2023f; Lee et al., 2024a). Methods like Distilling Step-by-Step (Hsieh et al., 2023) use teacher-generated rationales to train smaller models, improving performance with fewer examples. LaMini-LM (Wu et al., 2024c) creates a diverse instruction dataset with GPT-3.5 Turbo responses, enabling robust performance in smaller models.
2.3.3. Quantization
Quantization reduces the storage and computational demands of LLMs by converting floating-point representations into lower-precision formats, significantly cutting both storage requirements and computational complexity. Existing methods fall into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). Figure 9 illustrates the two quantization methods. Post-Training Quantization, applied after training, simplifies model compression without altering the architecture or requiring retraining, though it may result in precision loss. Consider a group or block of weights $\mathbf{w}$; the linear operation can be expressed as $y=\mathbf{w}\mathbf{x}$, while the quantized version is given by $y=Q(\mathbf{w})\mathbf{x}$. Generally, the quantization function $Q$ is defined as (Lin et al., 2024c): $Q(\mathbf{w})=\Delta\cdot\text{Round}\left(\frac{\mathbf{w}}{\Delta}\right),\quad\Delta=\frac{\text{max}(|\mathbf{w}|)}{2^{N-1}},$ where $N$ is the number of quantization bits, and $\Delta$ is the quantization scale factor determined by the absolute maximum value of $\mathbf{w}$. Quantization-Aware Training (QAT) enhances LLM efficiency by directly incorporating quantization into the training process, often resulting in higher accuracy than PTQ. During QAT, the forward pass utilizes quantized weights $Q(\mathbf{W})$ and activations $Q(\mathbf{X})$, while retaining full-precision values during the backward pass and for updating gradients to ensure stable learning dynamics. The comparisons of the post-training quantization methods are summarized in Table 1, detailing precision, addressed problems, and technical contributions of each method.
@@ 2.3.3. Quantization
Methods Bit Type Technical Contribution Problems
@@
OneBit (Xu et al., 2024b) 1-bit PTQ Quantization-aware knowledge distillation 1-bit quantization
@@ 2.3.4. Low-Rank Techniques
Low-rank techniques compress LLMs by approximating a high-dimensional weight matrix with two lower-dimensional matrices, reducing computational and memory requirements. A matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ is approximated as $\mathbf{W}\approx\mathbf{A}\times\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{m\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times n}$, with $r$ much smaller than $m$ or $n$, reducing the number of parameters. Building on this concept, Ji et al. (2024) propose a low-rank method tailored for LLMs, leveraging the observation that while LLMs have high-rank weights, their feature interactions tend to exhibit low-rank properties. The method estimates feature distributions using pooled covariance matrices and allocates distinct compression ratios to layers based on their sensitivity to low-rank compression. A Bayesian optimization strategy, using a Gaussian process as the surrogate model, optimizes the allocation of low-rank dimensions, ensuring the model maintains performance while achieving significant compression. Transitioning from model compression to fine-tuning, Cho et al. (2024) tackles system and data heterogeneity with the HETLORA method, which uses heterogeneous low-rank approximations to accommodate the diverse capabilities of clients and data complexities. By combining local rank self-pruning with sparsity-weighted aggregation, it balances high and low-rank LoRA modules, improving convergence speed and performance compared to uniform approaches. LLM-Neo (Yang et al., 2024g) combines knowledge distillation with low-rank adaptation (LoRA) to improve the efficiency of transferring knowledge from a teacher LLM to a compact student model.
@@ 3. Advanced Enhancement Strategies for Small Language Models
Training from Scratch MindLLM (Yang et al., 2023a) Bilingual models with advanced features.
MobiLlama (Thawakar et al., 2024) On-device SLM with dual objectives for efficiency and capability.
MobileLLM (Liu et al., 2024h) Optimizes LLM deployment on mobile with advanced architecture.
Supervised Fine-tuning MobileBERT (Sun et al., 2020) Compact BERT for efficient fine-tuning.
Alpaca 7B (Taori et al., 2023) Uses ChatGPT-generated tasks to tune Llama 7B.
RLHF (Ouyang et al., 2022) Trains using human-preferred data and reinforcement learning.
DPO (Rafailov et al., 2024) Dynamically adjusts log probabilities to prevent model degradation.
Data Quality in KD TinyStory (Eldan and Li, 2023) Enhances narrative coherence in child-friendly datasets.
AS-ES (Xi et al., 2024) Improves CoT by categorizing reasoning steps.
Self-Amplify (Bhan et al., 2024) Automates CoT data annotation for small models.
Distillation for SLM GKD (Agarwal et al., 2024) Aligns training and inference distributions using on-policy sequences.
DistiLLM (Ko et al., 2024) Uses skew KL divergence and adaptive off-policy for output utilization.
Adapt-and-Distill (Yao et al., 2021) Domain adapts both teacher and student models before distillation.
Quantization SmoothQuant (Xiao et al., 2023) Balances quantization difficulty using per-channel scaling.
BiLLM (Huang et al., 2024b) Applies Hessian-based metrics for binary residual approximation.
LLM-QAT (Liu et al., 2023a) Uses data-free knowledge distillation and logit distillation for fine-tuning.
PB-LLM (Shang et al., 2023) Binarizes non-salient weights while preserving others in higher precision.
OneBit (Xu et al., 2024b) Achieves near 1-bit quantization with minimal performance loss.
BitNet (Wang et al., 2023e) Introduces 1-bit Transformer architecture with BitLinear layers.
BitNet b1.58 (Ma et al., 2024) Implements a ternary weight system in enhanced BitNet.
LLM techniques for SLM Ma et al. (2023a) Combines filtering and re-ranking to improve Information Extraction tasks.
MoQE (Kim et al., 2023a) Applies quantization to expert weights to outperform dense models.
SLM-RAG (Liu et al., 2024g) Shows that SLMs with RAG can match LLM performance.
@@ 3.1. Innovative Training Methods for Small Language Models from Scratch
Architecture Design for SLMs When designing SLM architectures, parameter-sharing techniques are employed to minimize space usage and reduce the model’s size. As shown in the first part of Figure 10, parameter sharing is achieved by two approaches: (i) a single Feed-Forward Network (FFN) module is shared by every transformer layer. As shown in Figure 10 (1) middle, FFN layer sharing/reusing can maintain a smaller size while still benefiting from the depth and complexity gained through repeated processing of input data. This technique is firstly applied in MobiLlama (Thawakar et al., 2024) which surpasses the performance of existing SLMs of comparable size. (ii) Entire transformer blocks are shared. As shown in Figure 10 (1) right, Transformer Block-wise Sharing is another parameter-sharing approach that maintains depth and complexity. There are different transformer block-wise sharing strategies such as repeating the transformer blocks all over again or repeating the immediate transformer block. This technique is applied in MobileLLMs (Liu et al., 2024h) which has 125M and 350M parameters. MobileLLMs demonstrate performance improvements of 2.7% and 4.3%, respectively, compared to previous models with equivalent parameters. Moreover, they exhibit accuracy comparable to LLaMa-2-7B on API call tasks, highlighting the capabilities of smaller models in mobile environments.
@@
Data Construction For SLMs, the emphasis on data quality surpasses that of quantity and diversity (Yang et al., 2023a). Experiments demonstrate that using a quality filtering approach to remove low-quality data can lead to improved performance in SLMs (Yang et al., 2023a). Unlike large models, which can handle diverse and large datasets, SLMs benefit more from cleaner, high-quality data probably due to their limited capacity against noise. Generally, data processing has several steps: (i) Remove HTML, CSS, JS, and non-text elements for clean text; (ii) Filter low text-to-content ratio web pages; (iii) Deduplicate using SimHash (Dasgupta et al., 2011; Sadowski and Levin, 2007); (iv) Exclude sensitive/offensive content with heuristics and token replacements; (v) Remove self-repeating phrases of advertisements to enhance dataset informativeness (Chen et al., 2024d; Yang et al., 2023a). These steps collectively ensure that training data has high-quality, informative texts. SLMs also significantly benefit from these techniques. For example, MindLLMs (Yang et al., 2023a), which are bilingual lightweight language models (available in 1.3B and 3B versions), adopt these data processing techniques and achieve improved capability acquisition.